# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v21

@@ 3.27 Miscellaneous Applications
Sandathara et al. [582] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [583] assists children to learn to speak without stuttering. Nethmi et al. [584] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. Ranasinghe et al. [585] trained a Sinhala image caption generator model on Google translated MS COCO [586] dataset and tested it on Google translated Flickr8K [587] and MS COCO datasets.
@@
Rajitha et al. [438] has proposed a statistical machine translation and transliteration method to align Sinhala and English documents. Rajitha et al. [588] used the data set that they introduced in their previous work [94] to prove that task specific supervised distance learning metrics outperform their unsupervised counterparts, for document alignment. Fernando et al. [125] have used pre-trained multilingual language models (PMLMs) to improve document and sentence alignment between Sinhala–English, Tamil–English, and Sinhala–Tamil language pairs.
@@
Kumari and Hettiarachchi [589] proposed an algorithm for Sinhala topic modelling based on LDA [225] and RAKE [590]Batawalaarachchi [591] proposed two methods using statistical features to select words to be included in Sinhala document titles. Arambewela et al. [592] proposed a Sinhala writing assistant tool utilising CNNS.  Jayaweera et al. [593] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [594]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [98] has published two Sinhala BERT models: SinBERT-large111111111https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small112112112https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo113113113https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [595] on the OSCAR [140, 141] dataset. Wickramasinghe and de Silva [134] used Sinhala-English dictionary data [133] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub114114114https://bit.ly/3t3SKu7. Liyanage et al. [596] have created a small Sinhala treebank of 100 sentences. Burchell et al. [132].show that they obtain better results for the LID task for most of the languages compared to NLLB [123]. However, for Sinhala the results by Burchell et al. [132] are the same as that of NLLB. Minixhofer et al. [597] have used Sinhala from OPUS100 [120] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [598] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [573] and mT5 [287, 573] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [599]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [600]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [601].
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [616, 617, 618, 619, 620] or foreigners [621, 186, 622]115115115Note that [622] is an extension of [186].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@ 5 Conclusion
At this point, a reader might think, there seems to be a significant number of implementations of NLP for Sinhala. Therefore, how can one justify listing Sinhala as a resource poor language? The important point which is missing in that assumption is that in the cases of almost all of the above listed implementations and findings, the only thing that is publicly available for a researcher is a set of research papers. The corpora, tools, algorithm, and anything else that were discovered through these research are either locked away as properties of individual research groups or worse lost to the time with crashed ancient servers, lost hard drives, and expired web hosts. This reason and probably academic/research rivalry have caused these separate research groups not to cite or build upon the works of each-other. In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig. 7 in Appendix B). This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP work-flow. In their analysis of low resourced languages, Ranathunga and de Silva [642] observed that only 11.43% of Sinhala NLP papers have released the relevant data sets. Further, according to them, code being released sits at 9.71% while tools being released sits at 5.71%. While the 11.43% figure of data release may induce a feeling of availability, Ranathunga and de Silva [642] further observes that it is only 1.14% that has been released in public repositories. Research publications promising access to data and code only to be found lacking later is a common academic shortfall according to Gabelica et al. [36] however, given that Sinhala NLP is already having minimal work done as a whole compared to some other languages, we simply cannot afford to lose any of the generated data or code. In conclusion, what can be said is that, even though there are islands of implementations done for Sinhala NLP, they are of very small scale and/or are usually not readily accessible for further use and research by other researchers. Thus, so far, sadly, Sinhala stays a resource poor language.
@@ Appendix B Author Meta Analysis
Figure 7: The Probability of research from an institution citing that of other institutions. Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext117117117https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up. With all those limitations in mind, we still can make a few Interesting observations. The institute with the highest number of publications, UCSC - University of Colombo seems to be getting the most citations from most sources. However, they themselves seem to almost exclusively cite their own papers (0.7543). The only exception is the smaller number of works they cite (0.2114) from the Department of CSE - University of Moratuwa, the institute with the second highest number of publications. This leaves them with a 0.0343 probability of citing anyone else. Comparatively, Department of CSE - University of Moratuwa seems to be more egalitarian in citing. They have a lower self-citation probability (0.5251) and a higher probability of citing UCSC - University of Colombo (0.3513). This results in a probability of citing others at 0.1236. Both the Johns Hopkins University and University of Edinburgh prefer to cite Google instead of work from Sri Lanka. Faculty of IT - University of Moratuwa prefers to cite UCSC - University of Colombo (0.2339) rather than Department of CSE - University of Moratuwa (0.0968) which is from the same parent institution.

### v22

@@ 3.29 Miscellaneous Applications
Ranasinghe et al. [622] trained a Sinhala image caption generator model on Google translated MS COCO [623] dataset and tested it on Google translated Flickr8K [624] and MS COCO datasets.
@@
Rajitha et al. [455] has proposed a statistical machine translation and transliteration method to align Sinhala and English documents. Rajitha et al. [625] used the data set that they introduced in their previous work [97] to prove that task-specific supervised distance learning metrics outperform their unsupervised counterparts, for document alignment. Fernando et al. [129] have used pre-trained multilingual language models (PMLMs) to improve document and sentence alignment between Sinhala–English, Tamil–English, and Sinhala–Tamil language pairs.
@@
Kumari and Hettiarachchi [626] proposed an algorithm for Sinhala topic modelling based on LDA [232] and RAKE [627]Batawalaarachchi [628] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [629] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [630] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [631] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [632] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [633] have created a small Sinhala treebank of 100 sentences. Burchell et al. [136] show that they obtain better results for the LID task for most of the languages compared to NLLB [127]. However, for Sinhala the results by Burchell et al. [136] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Minixhofer et al. [634] have used Sinhala from OPUS100 [124] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [635] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [613] and mT5 [299, 613] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [636]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [637]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [638]. Senarathne [639] introduced a model to translate Sinhala pseudocode to computer code written in C#. Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [597] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [640] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [655, 656, 657, 658, 659] or foreigners [660, 193, 661]125125125Note that [661] is an extension of [193].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@ 5 Conclusion
At this point, a reader might think, there seems to be a significant number of implementations of NLP for Sinhala. Therefore, how can one justify listing Sinhala as a resource poor language? The important point which is missing in that assumption is that in the cases of almost all of the above listed implementations and findings, the only thing that is publicly available for a researcher is a set of research papers. The corpora, tools, algorithm, and anything else that were discovered through these research are either locked away as properties of individual research groups or worse lost to the time with crashed ancient servers, lost hard drives, and expired web hosts. This reason and probably academic/research rivalry has caused these separate research groups not to cite or build upon the works of each other. In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource-rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig. 7 in Appendix B). This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP workflow. In their analysis of low-resourced languages, Ranathunga and de Silva [25] observed that only 11.43% of Sinhala NLP papers have released the relevant data sets. Further, according to them, code being released sits at 9.71% while tools being released sits at 5.71%. While the 11.43% figure of data release may induce a feeling of availability, Ranathunga and de Silva [25] further observes that it is only 1.14% has been released in public repositories. Research publications promising access to data and code only to be found lacking later is a common academic shortfall according to Gabelica et al. [39] however, given that Sinhala NLP is already having minimal work done as a whole compared to some other languages, we simply cannot afford to lose any of the generated data or code. In conclusion, what can be said is that, even though there are islands of implementations done for Sinhala NLP, they are of very small scale and/or are usually not readily accessible for further use and research by other researchers. Thus, so far, sadly, Sinhala stays a resource-poor language.
@@ Appendix B Author Meta Analysis
Figure 7: The Probability of research from an institution citing that of other institutions. Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext127127127https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up. With all those limitations in mind, we still can make a few Interesting observations. The institute with the highest number of publications, UCSC - University of Colombo seems to be getting the most citations from most sources. However, they themselves seem to almost exclusively cite their own papers (0.7543). The only exception is the smaller number of works they cite (0.2114) from the Department of CSE - University of Moratuwa, the institute with the second highest number of publications. This leaves them with a 0.0343 probability of citing anyone else. Comparatively, Department of CSE - University of Moratuwa seems to be more egalitarian in citing. They have a lower self-citation probability (0.5251) and a higher probability of citing UCSC - University of Colombo (0.3513). This results in a probability of citing others at 0.1236. Both the Johns Hopkins University and University of Edinburgh prefer to cite Google instead of work from Sri Lanka. Faculty of IT - University of Moratuwa prefers to cite UCSC - University of Colombo (0.2339) rather than Department of CSE - University of Moratuwa (0.0968) which is from the same parent institution.